OpenStudy (anonymous):

Find the derivative of f(x) = 6 divided by x at x = -2.

OpenStudy (danjs):

\[6*\frac{ d }{ dx }\frac{ 1 }{ x }\]

OpenStudy (anonymous):

What is the d?

OpenStudy (fibonaccichick666):

that is just how you represent a derivative

OpenStudy (anonymous):

Would I just place in 2 for x?

OpenStudy (danjs):

6 times derivative of x^(-1)

OpenStudy (danjs):

-6x^-2

OpenStudy (fibonaccichick666):

you could also write \(=6(\frac{1}{x})'\)

OpenStudy (anonymous):

then where do I go from there?

OpenStudy (danjs):

put in x = -2

OpenStudy (fibonaccichick666):

but you do need to find the derivative of the section in parentheses first as dan has done

OpenStudy (danjs):

-6 / (-2)^2

OpenStudy (anonymous):

So 9?

OpenStudy (princeharryyy):

-6/x*2 = -6/4 = -3/2

OpenStudy (danjs):

-6 / 4

OpenStudy (anonymous):

OH okay I thought to divide then square

OpenStudy (princeharryyy):

no problem it was x^2

OpenStudy (fibonaccichick666):

order of operations there hun

OpenStudy (anonymous):

so -3/2

OpenStudy (danjs):

Good luck in calc!

OpenStudy (anonymous):

Thank you!!!!!